OpenStudy (anonymous):

Graph y = log5x and its inverse.

OpenStudy (anonymous):

@SinginDaCalc2Blues

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (campbell_st):

so do you know what the graph of \[y = \log_{5}(x)\] looks likes...?

OpenStudy (campbell_st):

do I have the right equation..?

OpenStudy (campbell_st):

because if thats the graph, we're in trouble...?

OpenStudy (anonymous):

yes ur equation is right

OpenStudy (anonymous):

which graph is the answer

OpenStudy (campbell_st):

ok... so start with a basic... do you know the equation of the inverse...?

OpenStudy (anonymous):

nop

OpenStudy (campbell_st):

ok... if I wrote \[x = \log_{5} (y)\] can you make y the subject...? which will give the inverse...

OpenStudy (anonymous):

so x^5=y

OpenStudy (campbell_st):

not quite y = 5^x... so if you use a table of values x = -2 -1 0 1 2 ------------------------ y = 0.04 0.2 1 5 25 do you have a graph that appears to go though these points...? if you do... the graph of \[y = \log_{5}(x)\] is reflected in the line y = x....

OpenStudy (anonymous):

did u see the choices that I post

OpenStudy (campbell_st):

so perhaps a really easy solution is to draw y = x on the 4 graphs... and see which has an original and reflection... that don't intersect..

OpenStudy (campbell_st):

yep... I saw them... trying to help you understand them

OpenStudy (anonymous):

yeah thnx

OpenStudy (campbell_st):

so a positive exponential graph looks like  maybe this helps